# Improved Tangential Interpolation-based Multi-input Multi-output Modal Analysis of a Full Aircraft

### v1

@@ Abstract
In the field of Structural Dynamics, modal analysis is the foundation of System Identification and vibration-based inspection. However, despite their widespread use, current state-of-the-art methods for extracting modal parameters from multi-input multi-output (MIMO) frequency domain data are still affected by many technical limitations. Mainly, they can be computationally cumbersome and/or negatively affected by close-in-frequency modes. The Loewner Framework (LF) was recently proposed to alleviate these problems with the limitation of working with single-input data only. This work proposes a computationally improved version of the Lowner Framework, or iLF, to extract modal parameters more efficiently. Also, the proposed implementation is extended in order to handle multi-input multi-output data in the frequency domain. This new implementation is compared to state-of-the-art methods such as the frequency domain implementations of the Least Square Complex Exponential method and the Numerical Algorithm for Subspace State Space System Identification on numerical and experimental datasets. More specifically, a finite element model of a 3D Euler-Bernoulli beam is used for the baseline comparison and the noise robustness verification of the proposed MIMO iLF algorithm. Then, an experimental dataset from MIMO ground vibration tests of a trainer jet aircraft with over 91 accelerometer channels is chosen for the algorithm validation on a real-life application. Excellent results are achieved in terms of accuracy, robustness to noise, and computational performance by the proposed improved MIMO method, both on the numerical and the experimental datasets. The MIMO iLF MATLAB implementation is shared in the work supplementary material.
@@
Loewner Framework , Multi-input multi-output , Tangential interpolation , Ground vibration testing , Modal analysis , Aeronautical structures , Vibration[label1]organization=Department of Aerospace Engineering, Universidad Carlos III de Madrid, addressline=Av.da de la Universidad 30, city=Leganés, postcode=28911, state=Madrid, country=Spain
@@
Pioneeringly, multi-input multi-output modal parameters extraction via the Loewner Framework is achieved;
This implementation of the Loewner Framework is more accurate and computationally efficient than its conventional counterpart, as well as other well-established methods ;
The proposed method is tested to identify the modal parameters of the full BAE Systems Hawk T1A aircraft benchmark.
@@ 1 Introduction
Modal analysis plays an important role during the design and validation of industrial systems and products [1]. In particular, the goal of EMA is to extract the modal parameters – natural frequencies (), damping ratios (), and mode shapes () – of the system under analysis [2]. These are generally employed for finite element model updating [3] or vibration-based Structural Health Monitoring (SHM) [4].
The modal parameters of the systems under scrutiny are extracted through system identification (SI) techniques, which are usually defined according to the type of data they are capable of processing. First, some techniques are meant for output-only data, where the input is unknown a priori, and others for input-output data. The former is the foundation of operational modal analysis and lies outside the scope of this work. The latter, known as experimental modal analysis (EMA), represents the context of this work. EMA methods can be further divided with respect to the number of input and output channels, considering one or more of them. Single-input single-output (SISO) and Single-input multi-output (SIMO) modal analysis are the most widely used methods for EMA as they are easier and cheaper to implement than their counterparts, especially multi-input multi-output (MIMO). That derives from the need of more advanced equipment and dedicated algorithms. In most cases, SIMO testing can suffice the scope of the test, such as when only one in-plane direction needs to be investigated [5]. However, at other times, a MIMO approach might be needed to fully characterise some of the modes of the system, such as in [6], where an out-of-plane mode is not fully characterised in the SIMO test.
In addition, despite SI being a mature and well-explored field [7], other-than-commercial openly available implementations of MIMO methods for modal analysis are very scarce. This is another hurdle for the implementation of more MIMO-based testing. Nevertheless, the MIMO algorithms have the same drawbacks as the existing SIMO methods: the ill-conditioning of the fitting matrix resulting in high computational cost and time [8]. Hence, a technique based on rational interpolation along tangential directions, the Loewner Framework (LF), has been extended for the extraction of the modal parameters from mechanical systems by the authors [9, 10, 11]. Hence, this work aims to introduce a novel computationally improved version of the LF (iLF) to extend the modal parameters extraction capability of the LF from SIMO to full MIMO capability. To achieve this this work tackles the following:
@@
again in Section 3, the newly-proposed iLF is compared to the standard LF approach, introduced in previous works [9], as well as to other state-of-the-art approaches - namely, the Least Square Complex Exponential (LSCE) [12] and the frequency domain numerical algorithm for subspace state space system identification (N4SID) [13]. Both accuracy and computational requirements are benchmarked for all procedures. The effects of artificially added white Gaussian noise are thoroughly investigated as well, to assess the effectiveness with increasingly noisy measurements;
@@
The proposed methodology is then experimentally validated in Section 4. Data from controlled laboratory GVT tests on a full-scale advanced training jet aircraft are used; both from a single wing and from the whole structure. As before, the results from other different classic algorithms are reported for direct comparability;
@@ 2.1 Multi-input multi-output testing for modal parameters extraction
As anticipated, MIMO testing is an advanced method in structural dynamics and modal analysis for experimentally extracting modal parameters from complex systems. Unlike traditional SISO and SIMO methods, MIMO testing involves simultaneously applying multiple inputs and measuring multiple responses. These multiple inputs might be in different positions on the specimen or excite the structure in different directions.
@@
111https://uk.mathworks.com/help/signal/ref/modalfit.html and n4sid
222https://uk.mathworks.com/help/ident/ref/n4sid.html, are used solely for benchmarking purposes.
A few examples of MIMO tests can be found in the current scientific literature with different applications. In [14], an inner disc and railway wheel rim are tested using multiple impact hammers as input. Notably, [15], while discussing the potential of digital twins for virtual testing, introduces the HYDRA test facility at the European Space Research and Technology Centre, which includes a 6 degrees-of-freedom vibration table for full-scale satellites. In aeronautics, [16] discusses the MIMO test, using two sting shakers as input, of a helicopter rotor spider and the advantages of the MIMO setup compared to a SISO setup for the characterisation of the system.
@@ 2.2 The Loewner Framework
The conventional LF algorithm has been applied to SI in electrical [17], aerodynamic [18], and, by the authors, to mechanical systems [9], but obtaining SIMO capability only. This Section first outlines the general background of the LF and then expands on the computational improvements brought by the iLF.
The LF fits Frequency Response Functions (FRFs) using rational interpolation, representing a function as the quotient of two polynomials. The Loewner matrix is constructed from pairs of complex numbers and as:
@@
For a linear time-invariant (LTI) dynamical system with inputs, outputs, and internal variables, the system is represented as:
where , , , and , , , and are the system matrices. Hence, the transfer function can be defined as follows:
The Loewner pencil, made up of the and the shifted Loewner matrix , satisfy the Sylvetser equation as such:
where , , and is the right interpolation data and , , and are the left interpolation data. Hence, for minimal data, where the right and left data have the same length:
where is the generalised tangential observability matrix and is the generalised tangential controllability matrix, resulting in:
The rational function is then:
For redundant data (the formulation used in this work), singular value decomposition (SVD) is used, and the realisation matrices are:
where and are, respectively, the generalised controllability and observability matrices of the realisation in Equation 3 for , as shown in [19].
Once, the modal parameters are then obtained from the system matrices. These insights represent the minimal theoretical background of the LF required for the understanding of this work, which aims at the extraction of modal parameters from mechanical MIMO systems. Nevertheless, the reader interested in a deeper insight into the LF background and its first applications in the field of electrical engineering is referred to [20, 19].
@@ 2.2.1 Computational improvements
As briefly mentioned in the earlier Sections, this work introduces an improved version of the LF. In this iLF implementation, the computational inefficiencies of the original implementation are amended to increase performance. Two main sources of inefficiencies are identified in the original LF implementation for the extraction of mechanical systems. The for
loops in the LF implementation originally reported in [17, 21, 22] and the way the matrices and are handled in the LF for the extraction of modal parameters extraction in [9, 23].
In the latter, the way the matrix and are obtained in the first author’s original contribution for modal parameters extraction has changed. This implementation created a descriptor state-space model with the dss
333https://uk.mathworks.com/help/control/ref/dss.html MATLAB function to obtain and , now this is dealt with matrix operations, more computationally efficient.
For the former, the most computationally cumbersome part was the coordinate transformation to obtain matrices with real entries from the LF. This iterative step in the implementation, once available in [22]444At the time of writing, it is no longer possible to download the original implementation from the original source, previously found at http://aci.rice.edu/system-identification/. An implementation, modified by the first author, is available at https://dspace.lib.cranfield.ac.uk/handle/1826/22465.. This is amended with a vectorised approach based on the Kronecker tensor product (kron
555https://uk.mathworks.com/help/control/ref/kron.html in MATLAB). These changes are predicted to have a noticeable effect on the computational performance. They are evaluated in the next Section.
Although the performance bottlenecks have been addressed, the iLF implementation, as described so far, is limited to the extraction of modal parameters from SIMO test settings due to the way the input data is handled. This is also amended to allow for the extraction of modal parameters from MIMO data.
@@ 3 Numerical study
In this Section, a 3D Euler Bernoulli beam is introduced to validate the newly proposed iLF MIMO implementation. The beam is modelled as a cantilever beam discretised in 8 elements and long 1.5 m (L). The beam material is Aluminium, the Young modulus (E) is 69 GPa and the density () is 2700 kgm-3. The beam cross-section is shown in Figure 1(b), alongside the beam discretisation (Figure 1(a)). As already mentioned earlier, this very simple FEM is intended as a controlled environment to precisely evaluate the capabilities of the proposed MIMO iLF approach; especially in terms of robustness to artificially-introduced measurement noise. Field test investigation on a real-life case study will follow in the next Section.
The excitation of the numerically simulated beam is applied at node # 2 (at a distance L/8 from the root) with a 1 N step input in the y and z directions. Rotations and x-axial displacement responses are ignored, resulting, therefore, in 16 non-zero displacement time series (in m; please note that the root node is constrained at the wall) and 2 input force time histories (in N, corresponding to the y- and z-direction input). All the input and the output signals are sampled at 800 Hz, allowing inspections up to 400 Hz due to the Nyquist criterion. A damping ratio of 3% is arbitrarily set for all modes. The MIMO Compliance FRFs () of the system take the shape of m:n:p
@@
The computed FRFs are then used to identify the modal parameters through the iLF MATLAB implementation, the standard LF, and the benchmark methods; N4SID and LSCE.
@@ 3.1 Modal parameters identification
The first step for assessing the suitability of iLF for the identification of modal parameters from MIMO measurements is to compare its performance. This is done here in terms of both precision and computational efficiency. Hence, the said FRFs were considered to identify the relevant modal parameters with N4SID, LSCE, LF (the standard implementation with the extra MIMO capability), and iLF to compare it with the analytical results from the 3D beam mass () and stiffness () matrices. The modal parameters are obtained by iterating over different orders for the four methods and extracting the relative modes using stabilisation diagrams between orders 32 (minimum order of the system) and 60. The resulting stabilisation diagrams are shown in Figure 2 and the identified are reported in Table 1.
From Figure 2 it is clear that all methods, LSCE apart, can identify 16 stable modes. The situation for the LSCE identification is not unusual. In fact, in [9], LSCE was found to struggle in situations where modes were close to each other. In the system under investigation, one can notice that modes #4-5 and 7-8 are relatively very close in frequency. On the computational side of things, it means that LSCE can only build rank-deficient matrices for approximating the system, hence misinterpreting it.
@@
As anticipated, LSCE cannot obtain any mode coherent with the analytical ones. In particular, the LSCE identified are overestimated by more than 400%. On the other hand, all other techniques showed similar results with an error, in terms of , of less than 0.05%. The same, even if not included in Table 1 for brevity, happens for , with the only exception of LSCE. For which errors over 29% are found.
In terms of , the Modal Assurance Criterion (MAC) value [24] between the modes identified and the analytical ones is 1, exception made for the LSCE identified , which have a MAC value of almost 0, showing total incompatibility.

### v2

@@ Abstract
In the field of Structural Dynamics, modal analysis is the foundation of System Identification and vibration-based inspection. However, despite their widespread use, current state-of-the-art methods for extracting modal parameters from multi-input multi-output (MIMO) frequency domain data are still affected by many technical limitations. Mainly, they can be computationally cumbersome and/or negatively affected by close-in-frequency modes. The Loewner Framework (LF) was recently proposed to alleviate these problems with the limitation of working with single-input data only. This work proposes a computationally improved version of the LF, or iLF, to extract modal parameters more efficiently. Also, the proposed implementation is extended in order to handle MIMO data in the frequency domain. This new implementation is compared to state-of-the-art methods such as the frequency domain implementations of the Least Square Complex Exponential method and the Numerical Algorithm for Subspace State Space System Identification on numerical and experimental datasets. More specifically, a finite element model of a 3D Euler-Bernoulli beam is used for the baseline comparison and the noise robustness verification of the proposed MIMO iLF algorithm. Then, an experimental dataset from MIMO ground vibration tests of a trainer jet aircraft with over 91 accelerometer channels is chosen for the algorithm validation on a real-life application. Its validation is carried out with known results from a single-input multi-output dataset of the starboard wing of the same aircraft. Excellent results are achieved in terms of accuracy, robustness to noise, and computational performance by the proposed improved MIMO method, both on the numerical and the experimental datasets. The MIMO iLF MATLAB implementation is shared in the work supplementary material.
@@
Loewner Framework , Multi-input multi-output , Tangential interpolation , Ground vibration testing , Modal analysis , Aeronautical structures , VibrationSee pages 1 of preprint.pdf
[label1]organization=Department of Aerospace Engineering, Universidad Carlos III de Madrid, addressline=Av.da de la Universidad 30, postcode=28911, city=Leganés, state=Madrid, country=Spain
@@
A computationally improved version of the Loewner Framework (iLF) is proposed;
Multi-input multi-output (MIMO) modal parameters extraction via the iLF is achieved;
The iLF is more accurate and computationally efficient than previous implementations;
The proposed method outperforms well-established methods, such as N4SID and LSCE;
The iLF is used for the modal identification of a BAE Systems Hawk T1A aircraft from MIMO data.
@@ 1 Introduction
Modal analysis plays an important role during the design and validation of industrial systems and products [1]. In particular, the goal of Experimental Modal Analysis (EMA) is to extract the modal parameters – natural frequencies (), damping ratios (), and mode shapes () – of the system under analysis [2]. These are generally employed for finite element model updating [3] or vibration-based Structural Health Monitoring (SHM) [4].
The modal parameters of the systems under scrutiny are extracted through System Identification (SI) techniques, which are usually defined according to the type of data they are capable of processing. First, some techniques are meant for output-only data, where the input is unknown a priori, and others for input-output data. The former is the foundation of operational modal analysis and lies outside the scope of this work. The latter, known as EMA, represents the context of this work. EMA methods can be further divided with respect to the number of input and output channels, considering one or more of them. Single-input single-output (SISO) and single-input multi-output (SIMO) modal analysis are the most widely used methods for EMA as they are easier and cheaper to implement than their counterparts, especially multi-input multi-output (MIMO). That derives from the need for more advanced equipment and dedicated algorithms. In most cases, SIMO testing can suffice the scope of the test, such as when only one in-plane direction needs to be investigated [5]. However, at other times, a MIMO approach might be needed to fully characterise some of the modes of the system, such as in [6], where an out-of-plane mode is not fully characterised in the SIMO test.
In addition, despite SI being a mature and well-explored field [7], other-than-commercial openly available implementations of MIMO methods for modal analysis are very scarce. This is another hurdle for the implementation of more MIMO-based testing. Nevertheless, the MIMO algorithms have the same drawbacks as the existing SIMO methods: the ill-conditioning of the fitting matrix resulting in high computational cost and time [8]. Hence, a technique based on rational interpolation along tangential directions, the Loewner Framework (LF), has been extended for the extraction of the modal parameters from mechanical systems by the authors [9, 10, 11].
Thus, this work aims to introduce a novel computationally improved version of the LF (iLF) to extend the modal parameters extraction capability of the LF from SIMO to full MIMO capability. The main innovative aspects are not limited to the extension from SIMO to MIMO datasets; as it will be discussed in further detail, the newly introduced version of the LF, iLF, is designed to be computationally more efficient than the traditional LF. This is a significant advancement, as the original LF suffered from computational bottlenecks, especially when handling large datasets or high-dimensional systems.
To achieve this aim, this work tackles the following:
@@
again in Section 3, the newly-proposed iLF is compared to the standard LF approach, introduced in previous works [9], as well as to other state-of-the-art approaches - namely, the Least Square Complex Exponential (LSCE) [12] and the frequency domain numerical algorithm for subspace state space system identification (N4SID) [13]. Both accuracy and computational requirements are benchmarked for all procedures. The effects of artificially added white Gaussian noise (WGN) are thoroughly investigated as well to assess the effectiveness with increasingly noisy measurements;
@@
The proposed methodology is then experimentally validated in Section 4. Data from controlled laboratory Ground Vibration Tests on a full-scale advanced training jet aircraft are used, both from a single wing and from the whole structure. As before, the results from other different classic algorithms are reported for direct comparability;
@@ 2.1 Multi-input multi-output testing for modal parameters extraction
As already hinted before, MIMO testing is an advanced method in structural dynamics and modal analysis for experimentally extracting modal parameters from complex systems. Unlike traditional SISO and SIMO methods, MIMO testing involves simultaneously applying multiple inputs and measuring multiple responses. These multiple inputs might be in different positions on the specimen or excite the structure in different directions.
@@
222https://uk.mathworks.com/help/signal/ref/modalfit.html and n4sid
333https://uk.mathworks.com/help/ident/ref/n4sid.html, are used solely for benchmarking purposes.
A few examples of MIMO tests can be found in the current scientific literature, with different applications. In [14], an inner disc and railway wheel rim are tested using multiple impact hammers as input. Notably, [15], while discussing the potential of digital twins for virtual testing, introduces the HYDRA test facility at the European Space Research and Technology Centre, which includes a 6 degrees-of-freedom vibration table for full-scale satellites. In aeronautics, [16] discusses the MIMO test, using two sting shakers as input, of a helicopter rotor spider and the advantages of the MIMO setup compared to a SISO setup for the characterisation of the system.
@@ 2.2 The Loewner Framework
The conventional LF algorithm has been applied to the modelling of a multi-port electrical system [17] and to aerodynamic model order reduction for aeroservoelastic modelling in [18]. The authors introduced LF for the identification of modal parameters from mechanical systems in [9], validated its computational performance in [10] and its noise robustness characteristic for SHM in [11]. However, these studies were limited in their applicability to SIMO systems only. This Section first outlines the general background of the LF and then expands on the computational improvements brought by the iLF and its extension to MIMO systems.
Lets begin by defining the Loewner matrix : Given a row array of pairs of complex numbers (,), ,…,, and a column array of pairs of complex numbers (,), ,…,, with , distinct, the associated , or divided-differences matrix is:
If there is a known underlying function , then and
Karl Löwner established a connection between and rational interpolation, also known as Cauchy interpolation [19]. This allows to define interpolants based on determinants of submatrices of . According to [20, 21], rational interpolants can be derived from . The approach based on the Loewner pencil is considered in this work. The Loewner pencil consists of the and matrices, where is the Shifted Loewner matrix, defined later.
@@
In order to describe how the LF works, let us consider a linear time-invariant dynamical system with inputs and outputs, and internal variables in descriptor-form representation, given by:
where is the internal variable, is the function’s input and is the output. The constant system matrices are:
a Laplace transfer function, , of can be defined in the form of a rational matrix function, when the matrix is non singular for a given finite value , such that :
Let us consider the more general case of tangential interpolation (rational interpolation along tangential directions [22]). The right interpolation data becomes:
Likewise, the left interpolation data:
and are the values at which is evaluated (the frequency bins in this application). The vectors and are, respectively, the right and left tangential general directions, randomly selected in practice [18], and and are the right and left tangential data. Linking and to the transfer function , associated with realisation in Equation 2, solves the rational interpolation problem:
such that the Loewner pencil satisfies Equation 7.
Now, let us consider a set of points in the complex plane, a rational function , such that , and let . Including the left and right data partitions, the following is obtained:
with . Thus, the matrix becomes:
Since and are scalars, the Sylvester equation is satisfied by as follows:
Now, Let us define the shifted Loewner matrix, , as the corresponding to :
Likewise, the Sylvester equation is satisfied as follows:
Let us focus on considering Equation 4 and matrix . As shown in [21], is set to per convention as its contributions are incorporated in the other matrices. So, Equation 4 becomes:
A realisation with the smallest possible dimension exists only if the system is fully controllable and observable. Therefore, assuming the data is sampled from a system whose transfer function is characterised by Equation 13, the generalised tangential observability, , and generalised tangential controllability, , are defined as follows [23]:
Now, let us incorporate Equations 14 and 15 into, respectively, Equations 9 and 11:
Firstly, let us consider the case with a minimal amount of data, where it is assumed that . This assumption is based on the fact that no duplicate data is permitted in and . Thus, rearranging Equations 14 and 15 into 16 and 17:
Then, letting the Loewner pencil be a regular pencil, in the sense of :
Accordingly, the interpolating rational function is defined as follows:
The derivation presented applies to the minimal data scenario, a case rarely encountered in real-world datasets. Nevertheless, the LF can be expanded to handle redundant data points effectively.
Firstly, let us assume:
Secondly, the short Singular Value Decomposition (SVD) of is computed:
where and . Thirdly, note that:
and likewise, such that and are, respectively, the generalised controllability and observability matrices for the system with . After verifying the right and left interpolation conditions, the Loewner realisation for redundant data is given by:
The formulation of Equation 24 — that is, the Loewner realisation for redundant data — will be employed throughout the remainder of this work. For a more detailed discussion of each step, the interested reader is referred to [21, 20]. Finally, the modal parameters of the system can then be extracted by eigenanalysis of the system matrices and in Equation 24.
@@ 2.2.1 Computational improvements
As briefly mentioned in the earlier Sections, this work introduces an improved version of the LF. In this iLF implementation, the computational inefficiencies of the original implementation are amended to increase performance and the implementation issues resolved to allow for MIMO modal identification capabilities.
In terms of inefficiencies in the original LF implementation for the extraction of mechanical systems, the following are identified:
1.
The for loops in the LF implementation originally reported in [17, 24, 25]: The most computationally cumbersome part was the coordinate transformation to obtain matrices with real entries from the LF. This was an iterative step in the LF original implementation, once available in [25]444At the time of writing, it is no longer possible to download the original implementation from the original source, previously found at http://aci.rice.edu/system-identification/. An implementation, modified by the first author, is available at https://dspace.lib.cranfield.ac.uk/handle/1826/22465.. This is amended in the iLF with a vectorised approach based on the Kronecker tensor product (
kron
555https://uk.mathworks.com/help/control/ref/kron.html in MATLAB). The code snippets of the former and current implementations are found in Appendix A.1. -
2.
The way the matrices and are handled in the LF for modal parameters extraction in [9, 26]: In the authors’ original contribution for modal parameters extraction with LF, a descriptor state-space model obtained from LF was converted to a continuous state-space model with the
dss
666https://uk.mathworks.com/help/control/ref/dss.html andss
777https://uk.mathworks.com/help/control/ref/ss.html MATLAB functions. This was a considerable bottleneck. Now amended with simple matrices operations to obtain the continuous state-space matrix for the extraction of modal parameters. See Appendix A.2 for the MATLAB code snippets of the original and improved implementation.
With respect to the capability expansion from SIMO to MIMO, the second improvement in the code allowed for a slick transformation also of multi-input data. With the old implementation, this was a cumbersome and computationally demanding process, which limited the LF to SIMO data. The iLF implementation is openly available in [27].
@@ 3 Numerical study
In this Section, a 3D Euler Bernoulli beam is introduced to validate the newly proposed iLF MIMO implementation. The beam is modelled as a cantilever beam discretised in 8 elements and 1.5 m long (L). The beam material is aluminium, with a Young modulus (E) of 69 GPa and a density () of 2700 kgm-3. The beam cross-section is shown in Figure 1(b), alongside the beam discretisation (Figure 1(a)). As already mentioned earlier, this very simple finite element model is intended as a controlled environment to evaluate the capabilities of the proposed MIMO iLF approach, especially in terms of robustness to artificially-introduced measurement noise. Field test investigation on a real-life case study will follow in the next Section.
The excitation of the numerically simulated beam is applied at node # 2 (at a distance L/8 from the root) with a 1 N step input in the y and z directions. Rotations and x-axial displacement responses are ignored, resulting, therefore, in 16 non-zero displacement time series (in m; please note that the root node is constrained at the wall) and 2 input force time histories (in N, corresponding to the y- and z-direction input). All the input and output signals are sampled at 800 Hz, allowing inspections up to 400 Hz due to the Nyquist criterion. A damping ratio of 3% is arbitrarily set for all modes. The MIMO Compliance FRFs () of the system take the shape of m:n:p
@@
The computed FRFs are then used to identify the modal parameters through the iLF, the standard LF, and the two benchmark methods – N4SID and LSCE. All these algorithms were implemented in MATLAB.
@@ 3.1 Modal parameters identification
The first step for assessing the suitability of iLF for the identification of modal parameters from MIMO measurements is to compare its performance. This is done here in terms of both precision and computational efficiency. Hence, the said FRFs were considered to identify the relevant modal parameters with N4SID, LSCE, LF (the standard implementation with the extra MIMO capability), and iLF to compare it with the analytical results from the 3D beam mass () and stiffness () matrices. The modal parameters are obtained by iterating over different orders for the four methods and extracting the relative modes using stabilisation diagrams between orders 32 (minimum order of the system) and 60. The resulting stabilisation diagrams are shown in Figure 2, and the identified are reported in Table 1.
From Figure 2, it is clear that all methods, LSCE apart, can identify 16 stable modes. The relatively poor performance of the LSCE algorithm is not unusual. In fact, in [9], LSCE was found to struggle in situations where modes were close to each other. In the system under investigation, one can notice that modes #4-5 and 7-8 are relatively close in frequency. On the computational side of things, it means that LSCE can only build rank-deficient matrices for approximating the system, hence misinterpreting it, or partially identifying it.
@@
As anticipated, LSCE cannot obtain the full set of modes of the system. In particular, the LSCE identified only the last four modes , with absolute errors orders of magnitude higher than the other methods (the maximum absolute error is found for mode 16 and was over 12%). On the other hand, all other techniques showed similar results with an error, in terms of , of less than 0.05%. The same, even if not included in Table 1 for brevity, happens for , with the only exception of LSCE. For which errors over 29% are found.
In terms of , the Modal Assurance Criterion (MAC) value [28] between the modes identified and the analytical ones is (close to) 1, exception made for the modes LSCE could not identify. Table 2 shows the diagonal terms (off-diagonal terms are negligible) of the MAC values between the analytical and those identified.
Mode | N4SID | LSCE | LF | iLF |
1 | 0.915 | - | 0.915 | 0.915 |
2 | 0.974 | - | 1 | 1 |
3 | 1 | - | 1 | 1 |
4 | 1 | - | 1 | 1 |
5 | 1 | - | 1 | 0.997 |
6 | 0.998 | - | 0.998 | 0.998 |
7 | 0.999 | - | 0.999 | 0.999 |
8 | 0.999 | - | 0.999 | 0.999 |
9 | 0.999 | - | 0.999 | 0.999 |
10 | 1 | - | 0.999 | 1 |
11 | 1 | - | 1 | 1 |
12 | 1 | - | 1 | 1 |
13 | 1 | 1 | 1 | 1 |
14 | 1 | 1 | 1 | 1 |
15 | 1 | 1 | 1 | 1 |
16 | 1 | 1 | 1 | 1 |